OpenStudy (anonymous):

given the following revenue and cost functions, find the equilibrium point to the nearest tenth. R(x)=59x-2x^2; C(x)=23x+95

OpenStudy (mathmate):

equilibrium point would be when R(x)=C(x), or \(59x-x^2=23x+95\) which simplifies to \(x^2-36x+95=0\) subject to x\(\ge\)0 however, in this case, you end up with two possible solutions.

OpenStudy (anonymous):

Thank you! Yes i think i got it then i would plug \[2x^2-36x+95\] into the quadratic formula to get 14.8 or 3.2 as my final answer

OpenStudy (mathmate):

Not too sure where you got the \(\color{red}{2}\) in \(\color{red}{2}x^2−36x+95=0\) That's where our answer differ.